Determine the solution of sec^2 theta -4=0 in the interval 0 b) if csc theta = 10/8 determine tan theta if pi/2
Answer by Alan3354(69443)

Determine the solution of sec^2 theta -4=0 in the interval 0 < theta < 2pi
sec^2 = 4
sec = ±2
cos = ±1/2
theta = pi/3, 2pi/3, 4pi/3, 5pi/3
============
b) if csc theta = 10/8 determine tan theta if pi/2 < theta < pi
In Q2
sin = 4/5
cos = 3/4
tan = sin/cos = -4/3 (negative in Q2)